OpenStudy (anonymous):

A) x : 2 4 6 y : 10 20 30 B) 5x+3y= -2 C) I don't know how to post the graph. :( @noneedtoknow

OpenStudy (campbell_st):

A) looks like the equaton y = 5x what does the question ask you to do...?

OpenStudy (anonymous):

yes, A) is y=5x. @Breezehh do you know how to prove that that is the correct answer/equation?

OpenStudy (anonymous):

@campbell_st "Find the rate of change for each table, equation, or graph. How did you find the rate of change?"

OpenStudy (anonymous):

@noneedtoknow No, I have no idea.

OpenStudy (anonymous):

Okay, so try inputting a set of the date into the equation y=5x

OpenStudy (anonymous):

So, Like. 10 = 5(2) ?

OpenStudy (anonymous):

yes!:) now solve it

OpenStudy (campbell_st):

ok... so you are asked to find the slope of the line basically in the 1st equation its y = 5x so the slope is 5 or rate of change in y with respect to x is 5 B) write it in slope intercept form subtract 5x from both sides 3y = -5x - 2 divide every term by 3 y = -5/3 x - 2/3 slope of the line or rate of change is -5/3 C) find 2 points on the line and find the slope, rise/run... this will be the rate of change

OpenStudy (anonymous):

Thank you both! I still have like, NO idea what I'm doing but this kind of clears it up!

OpenStudy (anonymous):

@noneedtoknow